# JailGuard: A Universal Detection Framework for LLM Prompt-based Attacks

### v3

@@ 7. Related work
LLM Attack And Defense Supplement to the prompt-based attack methods in § 2, researchers proposed other methods to automatically generate jailbreak and hijacking prompts (Niu et al., 2024; Geiping et al., 2024; Deng et al., 2023b; Liu et al., 2024b; Zhang et al., 2024; Andriushchenko et al., 2024; Deng et al., 2023a; Shen et al., 2024). Geiping et al. (Geiping et al., 2024) construct misleading, misinformation, and other non-jailbreaking attack instructions based on the existing jailbreak attacks (Zou et al., 2023). Unfortunately, we cannot find available open-source code or datasets of their attacks. Researchers also pay attention to other aspects of LLM security, e.g., backdoor attack (Abdelnabi et al., 2023; Xu et al., 2023; Huang et al., 2023b), privacy stealing attack (Sha and Zhang, 2024). We focus on the defense of multi-modal prompt-based attacks in this paper, which uses prompts as the carrier and does not require finetuning or modification of the target LLMs and systems.
@@
To defend against LLM attacks, in addition to the baselines in § 6.1, researchers have proposed other methods (Kumar et al., 2023; Chen et al., 2024b; Pi et al., 2024; Zhou et al., 2024; Wu et al., 2024; Zhou et al., 2024). Kumar et al. (Kumar et al., 2023) designed a detection method that splices the input text and applies a safety filter on all substrings to identify toxic content. However, this method will have a significant overhead on a long input prompt. Similar to Self-reminder (Xie et al., 2023), Self-defend (Wu et al., 2024) uses system prompts to ask LLMs to self-check whether the given input is an attack input. Unfortunately, we cannot find available open-source implementation. In this paper, we compare JailGuard with 12 state-of-the-art open-sourced detection and defense methods.
@@
Adversarial Attack and Defense in DNNs. White box attacks assume the attacker has full knowledge of the target model, including its architecture, weights, and hyperparameters. This allows the attacker to generate adversarial examples with high fidelity using gradient-based optimization techniques, such as FGSM (Goodfellow et al., 2014), BIM (Kurakin et al., 2018), PGD (Madry et al., 2017), Square Attack (Andriushchenko et al., 2020). AutoAttack (Croce and Hein, 2020) has been proposed as a more comprehensive evaluation framework for adversarial attacks. Recently, researchers have also been exploring the use of naturally occurring degradations as forms of attack perturbations. These include environmental and processing effects like motion blur, vignetting, rain streaks, varying exposure levels, and watermarks (Gao et al., 2022; Guo et al., 2020; Jia et al., 2020; Tian et al., 2021; Hou et al., 2023).
Adversarial defense can be categorized into two main types: adversarial training and adversarial purification (Nie et al., 2022). Adversarial training involves incorporating adversarial samples during the training process (Goodfellow et al., 2014; Madry et al., 2017; Athalye et al., 2018; Rade and Moosavi-Dezfooli, 2021; Ding et al., 2018), and training with additional data generated by generative models (Sehwag et al., 2021). On the other hand, adversarial purification functions as a separate defense module during inference and does not require additional training time for the classifier (Guo et al., 2017; Xu et al., 2017; Sun et al., 2019; Ho and Vasconcelos, 2022).
@@ 8. Discussion
JailGuard EnhancementJailGuard requires several additional LLM queries to generate variant responses and detect attacks. Even if it can generate a smaller number of variants (i.e., $N=4$), this extra runtime overhead is still unavoidable. Moreover, fewer variants lead to a degradation in detection results. Developing a more effective mutation strategy that maintains high detection accuracy with a lower query budget is a critical area for future research. One possible solution is to utilize speculative small models (e.g., GPT-3) to generate variant responses, thus reducing queries costs. ❷ The heuristic benign instructions are the main causes of the false positives in JailGuard (e.g., ‘Suggest some names for a friendly group in telegram’). They have no clear answers and their responses are prone to high divergence, which significantly contributes to false positives in detection. Identifying such heuristic benign questions and mitigating false positives in attack detection is a crucial challenge for enhancing JailGuard. One potential approach involves designing an AI-based filter to automatically filter out these heuristic benign inputs. ❸ JailGuard currently implements 18 mutators and a set of mutator combination policies for the inputs on text and image modalities. With the development of MLLMs, audio input is becoming another important modality (e.g., GPT-4o (OpenAI, 2024)). Existing work (Wu et al., 2023; Abdullah et al., 2021) has pointed out the characteristics of poor robustness and transferability of audio adversarial attacks. Based on such observations, although there are currently no relevant MLLM audio attack methods and datasets, the detection metric in JailGuard (i.e., divergence) still has feasibility in detecting audio attacks. How to design mutators for audio attacks will be a potential future direction. Our framework and dataset will be continuously updated and any new appearing attacks will be further collected and evaluated. You can find the latest information on our website (our, 2024).
@@
Diverse LLM Attacks Detection ❶ As an emerging research field, the security of large models has received widespread attention from researchers and industry. It is significant to add more types of attack inputs (e.g., data poisoning (Yan et al., 2023) and backdoor (Abdelnabi et al., 2023; Xu et al., 2023), and misinformation (Geiping et al., 2024) ) and build a comprehensive and universal benchmark for LLM defense. ❷ Our detection method fundamentally leverages the inherent non-robustness of attacks. Consequently, the vulnerabilities introduced by data poisoning and model backdoors, which also exhibit this non-robustness, could potentially be identified by our detection framework. A crucial future direction involves designing defense methods that are both effective and efficient, capable of generalizing across various types of attack inputs. Successfully achieving this would significantly enhance the deployment and application of trustworthy Language Model (LM) systems, contributing to their overall reliability and security.
@@ 9. Conclusion
In this paper, we propose JailGuard, a universal detection framework that detects both jailbreaking and hijacking attacks for LLM systems on both image and text modalities. To comprehensively evaluate the detection effect of JailGuard, we construct the first comprehensive LLM prompt-based attack dataset, covering 15 jailbreaking and hijacking attacks and 11,000 items of data on image and text modalities. Our experiment results show that JailGuard achieves the best detection accuracy of 86.14%/ 82.90% on text/image inputs, significantly outperforming state-of-the-art defense methods by 11.81%-25.73% and 12.20%-21.40%.

### v4

@@ 7. Related work
LLM Attack and Defense. Supplement to the prompt-based attack methods in § 2, researchers proposed other methods to automatically generate jailbreak and hijacking prompts (Niu et al., 2024; Geiping et al., 2024; Deng et al., 2023b; Liu et al., 2024d; Zhang et al., 2024; Andriushchenko et al., 2024; Deng et al., 2023a; Shen et al., 2024b). Geiping et al. (Geiping et al., 2024) construct misleading, misinformation, and other non-jailbreaking attack instructions based on the existing jailbreak attacks (Zou et al., 2023). Unfortunately, we cannot find available open-source code or datasets of their attacks. Researchers also pay attention to other aspects of LLM security, e.g., backdoor attack (Abdelnabi et al., 2023; Xu et al., 2023; Huang et al., 2023b), privacy stealing attack (Sha and Zhang, 2024). In this paper, we focus on the defense of multi-modal prompt-based attacks, which use prompts as the carrier and do not require finetuning or modification of the target LLMs and systems.
@@
To defend against LLM attacks, in addition to the baselines in § 6.1, researchers have proposed other methods (Kumar et al., 2023a; Chen et al., 2024b; Pi et al., 2024; Zhou et al., 2024; Wu et al., 2024; Zhou et al., 2024). Kumar et al. (Kumar et al., 2023a) designed a detection method that splices the input text and applies a safety filter on all substrings to identify toxic content. However, this method will have a significant overhead on a long input prompt. Similar to Self-reminder (Xie et al., 2023), Self-defend (Wu et al., 2024) uses system prompts to ask LLMs to self-check whether the given input is an attack input. Unfortunately, we cannot find available open-source implementation. In this paper, we compare JailGuard with 12 state-of-the-art open-sourced detection and defense methods.
@@
Adversarial Attack and Defense in DNNs. White box attacks assume the attacker has full knowledge of the target model, including its architecture, weights, and hyperparameters. This allows the attacker to generate adversarial examples with high fidelity using gradient-based optimization techniques, such as FGSM (Goodfellow et al., 2014), BIM (Kurakin et al., 2018), PGD (Madry et al., 2017), Square Attack (Andriushchenko et al., 2020). AutoAttack (Croce and Hein, 2020) has been proposed as a more comprehensive evaluation framework for adversarial attacks. Recently, researchers have also been exploring the use of naturally occurring degradations as forms of attack perturbations. These include environmental and processing effects like motion blur, vignetting, rain streaks, varying exposure levels, and watermarks (Gao et al., 2022; Guo et al., 2020; Jia et al., 2020; Tian et al., 2021; Hou et al., 2023). Adversarial defense can be categorized into two main types: adversarial training and adversarial purification (Nie et al., 2022). Adversarial training involves incorporating adversarial samples during the training process (Goodfellow et al., 2014; Madry et al., 2017; Athalye et al., 2018; Rade and Moosavi-Dezfooli, 2021; Ding et al., 2018), and training with additional data generated by generative models (Sehwag et al., 2021). On the other hand, adversarial purification functions as a separate defense module during inference and does not require additional training time for the classifier (Guo et al., 2017; Xu et al., 2017; Sun et al., 2019; Ho and Vasconcelos, 2022).
Randomized Data Smoothing. Researchers have proposed randomized smoothing to provide certified adversarial robustness (Cohen et al., 2019; Fischer et al., 2020; Hao et al., 2022; Ye et al., 2020; Zhao et al., 2022; Zeng et al., 2023). Randomized smoothing constructs multiple copies of the original input and then perturbs them by introducing Gaussian noise (Cohen et al., 2019), rotating images (Fischer et al., 2020), masking texts (Zeng et al., 2023), etc. Finally, it ensembles and aggregates the model’s outputs to these perturbated copies and selects the major class of these outputs as the final output. Based on the concept of randomized smoothing, SmoothLLM (Robey et al., 2023) first duplicates and perturbs copies of the given input and then uses refusal keywords to distinguish blocked attack responses from normal responses and aggregates them to obtain the final LLM response.
In this paper, we propose JailGuard that utilizes the differences between LLM responses to input variants to detect various prompt-based attacks. The mutators implemented in JailGuard can also be classified as part of the broad randomized smoothing framework, which encompasses various input noise-based methods. However, JailGuard’s methodology differs fundamentally from traditional randomized smoothing techniques (Cohen et al., 2019; Robey et al., 2023) that aggregate outputs of multiple copies. JailGuard analyzes the divergence patterns in LLM responses to detect potential attacks. Such a distinct approach sets JailGuard apart within the randomized smoothing instances. In addition to random mutators inspired by existing work (Cohen et al., 2019; Zeng et al., 2023; Cubuk et al., 2020; Bai et al., 2022), JailGuard further proposes semantic-guided mutators and the mutator combination policy. Ultimately, in the experiments, JailGuard achieves better detection results than baselines (including SmoothLLM).
@@ 8. Discussion
Alternative Solutions in JailGuard.Other Embedding models. JailGuard uses the embedding model ‘en_core_web_md’ to convert LLM responses into response vectors. We have tried other embedding models as alternative solutions, such as ‘en_core_web_lg’ model from the ‘spaCy’ library, and ‘bert-base-uncased’ model from ‘google-bert’ community (using the mean of the last layer embeddings as the response vector). Our experiment on 1,000 samples from the complete text dataset shows that the detection effects of using different embedding models are very close. Specifically, the average accuracy of the ‘en_core_web_md’ model on 8 single text mutators is 81.70%, and the average accuracy of separately using ‘en_core_web_lg’ and ‘bert-base-uncased’ are 81.90% and 80.20%, with an accuracy change of less than 2%. Considering that the sizes of these two alternative models are larger than ‘en_core_web_md’ and they introduce larger memory and time overhead while converting response vectors, JailGuard uses the ‘en_core_web_md’ model as the default setting. ❷ Mean Square Error (MSE). In addition to KL divergence, there is another alternative solution to measure the differences between variant responses, i.e., directly calculate MSE between the rows of the similarity matrix $S$ from Equation 4, and distinguish attack samples from benign samples based on the values of MSE. We can have a $N\times N$ MSE matrix $D^{m}$ and each element $D^{m}_{i,j}$ can be calculated as $D^{m}_{i,j}=\frac{1}{N}\sum_{k=1}^{N}\left(S_{i,k}-S_{j,k}\right)^{2}$. We randomly select a subset with 1,000 samples from the text dataset and conduct comparative experiments on the text mutators. The experimental results show that the average accuracy on the text mutators is 78.20% with $\theta^{m}=0.1$ obtained from the training set, which is marginally lower than the 80.40% accuracy achieved using KL divergence. Our analysis shows that the MSE distributions of benign samples and attack samples are relatively close, therefore, the detection accuracy of using MSE is sensitive to threshold selection. Applying a threshold $\theta^{m}$ obtained from the training set may lead to false positives and lower accuracy on the test set. Considering the detection effect, JailGuard finally adopts KL divergence as the default solution for attack detection. However, it is worth noting that the MSE method offers computational simplicity and achieves comparable detection performance with KL divergence. This makes it a viable alternative solution for resource-constrained environments.
@@
JailGuard Enhancement.JailGuard requires several additional LLM queries to generate variant responses and detect attacks. Even if it can generate a smaller number of variants (i.e., $N=4$), this extra runtime overhead is still unavoidable. Moreover, fewer variants lead to a degradation in detection results. Developing a more effective mutation strategy that maintains high detection accuracy with a lower query budget is a critical area for future research. One possible solution is to utilize small models (e.g., GPT-3) to perform speculative decoding and generate variant responses, thus reducing query costs. However, the lack of safety alignment or capability in speculative models may lead to a degradation in detection results. How to find suitable speculative models and enhance the detection process to achieve better detection performance will be a future direction. ❷ The heuristic benign instructions are the main causes of the false positives in JailGuard (e.g., ‘Suggest some names for a friendly group in telegram’). They have no clear answers and their responses are prone to high divergence, which significantly contributes to false positives in detection. Identifying such heuristic benign questions and mitigating false positives in attack detection is a crucial challenge for enhancing JailGuard. In addition, we observe that in the experiment, seemingly toxic benign prompts can easily cause false positives in existing detection methods and JailGuard. The safety alignment mechanism of LLM has a certain probability of providing refusal responses to seemingly harmful prompts. How to improve JailGuard and avoid such false positives is also a future research direction. One potential approach involves designing an AI-based filter to automatically filter out these heuristic or seemingly toxic benign inputs. ❸ JailGuard currently implements 18 mutators and a set of mutator combination policies for the inputs on text and image modalities. With the development of MLLMs, audio input is becoming another important modality (e.g., GPT-4o (OpenAI, 2024)). Existing work (Wu et al., 2023; Abdullah et al., 2021) has pointed out the characteristics of poor robustness and transferability of audio adversarial attacks. Based on such observations, although there are currently no relevant MLLM audio attack methods and datasets, the detection metric in JailGuard (i.e., divergence) still has feasibility in detecting audio attacks. How to design mutators for audio attacks will be a potential future direction. ❹ JailGuard has currently collected 15 prompt-based attack methods targeting LLM and MLLM and has built a dataset containing 11,000 items of data whose scale significantly exceeds the dataset used in existing detection baselines (Robey et al., 2023; Xie et al., 2023). Although our experiments show that JailGuard achieves better detection results than baselines, its detection effect may not be maintained on some unseen attack methods. How to update the dataset and JailGuard and continuously extend them on various representative new attack methods will be our future work. Our framework and dataset will be continuously updated and any new appearing attacks will be further collected and evaluated. You can find the latest information on our website (our, 2024).
@@
Diverse LLM Attacks Detection. ❶ As an emerging research field, the security of LLM systems has received widespread attention from researchers and industry. It is significant to add more types of attack inputs (e.g., data poisoning (Yan et al., 2023) and backdoor (Abdelnabi et al., 2023; Xu et al., 2023), and misinformation (Geiping et al., 2024) ) and build a comprehensive and universal benchmark for LLM defense. ❷ Our detection method fundamentally leverages the inherent non-robustness of attacks. Consequently, the vulnerabilities introduced by data poisoning and model backdoors, which also exhibit this non-robustness, could potentially be identified by our detection framework. A crucial future direction involves designing defense methods that are both effective and efficient, capable of generalizing across various types of attack inputs. Successfully achieving this would significantly enhance the deployment and application of trustworthy Language Model (LM) systems, contributing to their overall reliability and security.
@@ 9. Conclusion
In this paper, we propose JailGuard, a universal detection framework that detects both jailbreaking and hijacking attacks for LLM systems on both image and text modalities. To comprehensively evaluate the detection effect of JailGuard, we construct the first comprehensive prompt-based attacks dataset, covering 15 jailbreaking and hijacking attacks on LLM systems and 11,000 items of data on image and text modalities. Our experiment results show that JailGuard achieves the best detection accuracy of 86.14%/ 82.90% on text/image inputs, significantly outperforming state-of-the-art defense methods by 11.81%-25.73% and 12.20%-21.40%.
Acknowledgements.
The authors would like to thank the anonymous reviewers for their insightful comments and valuable suggestions. This work is supported partially by the National Key Research and Development Program of China (2023YFB3107400), the National Natural Science Foundation of China (62006181, 62132011, 62161160337, 62206217, U20A20177, U21B2018), and the Shaanxi Province Key Industry Innovation Program (2021ZDLGY01-02 and 2023-ZDLGY-38). Thanks to the New Cornerstone Science Foundation and the Xplorer Prize. This research is supported by the National Research Foundation, Singapore, the Cyber Security Agency under its National Cybersecurity R&D Programme (NCRP25-P04-TAICeN), and DSO National Laboratories under the AI Singapore Programme (AISG2-GC-2023-008). It is also supported by the National Research Foundation, Prime Minister’s Office, Singapore under the Campus for Research Excellence and Technological Enterprise (CREATE) programme.